# A Bayesian Modeling Framework for Estimation and Ground Segmentation of Cluttered Staircases

### v1

@@ VI Conclusion
In this paper we presented a novel approach to staircase estimation in cluttered real-world environments. We introduced a split staircase state-space model, and a Bayesian inference framework utilizing this model for robust state estimation. Our method effectively addresses challenges posed by noise, clutter, and occlusions, leading to significant improvements in accuracy compared to existing approaches. We demonstrated the robustness and effectiveness of our approach through extensive experiments on staircases captured both on real robots and in simulation, showcasing its ability to accurately estimate staircase parameters and location even in complex scenarios. Moreover, our method achieves real-time performance, making it suitable for deployment on robotic systems operating in dynamic environments.
@@
In the future, we aim to utilize the segmented stair surface regions for quadruped motion planning, enabling safe navigation on cluttered staircases. Additionally, we will investigate approaches to enable fusion across multiple robots using methods such as Covariance Intersection for deployment on a multi-robot system [sriganesh2024systemdesign] in multi-staircase environments.
[…]

### v2

@@ VI Conclusion
In this paper we presented a novel approach to staircase estimation in cluttered real-world environments. We introduced a split staircase state-space model, and a Bayesian inference framework utilizing this model for robust state estimation. Our method effectively addresses challenges posed by noise, clutter, and occlusions, leading to significant improvements in accuracy compared to existing approaches. We demonstrate the robustness and effectiveness of our approach through extensive experiments on staircases captured both on real robots and in simulation, showcasing its ability to accurately estimate staircase parameters and location even in complex scenarios. Moreover, our method achieves real-time performance, making it suitable for deployment on robotic systems operating in dynamic environments.
@@
In the future, we aim to utilize the segmented stair surface regions for quadruped motion planning, enabling safe navigation on cluttered staircases. Additionally, we will investigate approaches to enable fusion across multiple robots using methods such as Covariance Intersection for deployment on a multi-robot system [21] in multi-staircase environments.